# Global-Decision-Focused Neural ODEs for Proactive Grid Resilience Management

### v2

@@ III-C Model Estimation
The learning of GDF Neural ODEs is carried out through stochastic gradient descent, where the gradient is calculated using a novel algorithm based on differentiable optimization techniques [28, 39, 40]. To enable differentiation through the $\arg\min$ operator in (2) embedded in the global-decision-focused loss, we relax the decision variable $\bm{x}$ from a potentially discrete space to a continuous space. For combinatorial optimization problems, the problem is reformulated as a differentiable quadratic program, and a small quadratic regularization term is added to ensure continuity and strong convexity [37]. Formally, we replace the original objective in (2) with the following:
@@
where $\rho>0$ ensures differentiability. More implementation details can be found in the supplementary material.
@@ IV Application: Mobile Generator Deployment
Figure 3: A synthetic example of the mobile generator deployment problem for a system with three cities and five generators ($Q_{w}=5$). The $y$-axis represents the number of outaged households. In this example, the uniform travel time $\delta_{t}=10$, transportation cost is set to $c=400$, the customer interruption cost is $\tau=1$, and the operational cost is $\gamma=2$. Four methods are compared on out‑of‑sample data: the proposed GDF framework (regret = $183.47$), a two‑stage approach (regret = $488.45$), an online baseline with observation lag of 5 (regret = $2846.13$), and the optimal solution with the groundtruth. Details of the synthetic data setup are provided in Section V-A.
Our method is broadly applicable to various proactive decision-making problems in grid operations, particularly in response to potential hazard events. This section highlights the adaptability of GDF by applying it to a mobile generator deployment task [16, 17, 18], a representative PATOG problem.
@@
Formally, let $Q_{w}\in\mathbb{Z}_{+}$ denote the initial inventory of generators at warehouse $w$. For simplicity, we assume that warehouses also function as staging areas, where generators are initially stored and returned after deployment. We assume that each generator has a fixed maximum capacity, capable of supplying electricity to $N_{g}$ customers, and that a single extreme event is anticipated. The planning horizon is discretized into uniform time periods $\mathcal{T}=\{1,2,\dots,T\}$, where ${T}$ represents the expected duration of the outage. Let $\mathcal{K}=\{1,2,\dots,K\}$ be the set of $K$ service units (e.g., cities or counties) where generators can be deployed, and $\mathcal{W}=\{1,\dots,W\}$ be the set of warehouses where all generators are initially stored. We denote the stock of generators in unit $k$ at time $t$ as $q_{tk}$.
@@
which consists of three components: ($i$) Transportation Cost: The cost of delivering generators from warehouses to service units. ($ii$) Operation Cost: The fixed operational cost per unit time ($\gamma$) for deployed generators. ($iii$) Outage Cost: Economic loss due to power outages, determined by the number of customers experiencing outages given the generator deployment decision $\bm{x}$. For simplicity, we define the outage cost as:
@@
where $\tau$ denotes unit customer interruption cost, an adapted version of the Value of Lost Load (VoLL) [46], representing the economic impact per household outage per day, and $Y_{k}(t)$ denotes the number of outages in service unit $k$ at time $t$, extracted from the state $\mathbf{S}$ according to (5). This framework primarily evaluates system functionality based on the number of customers experiencing outages ($Y$). It is worth emphasizing that it is flexible to incorporate additional metrics, including the number of recovered customers ($R$) and unaffected customers ($U$), if needed.
The transportation and inventory constraints of the mobile genereator deployment are as follows:
@@ V-A Dataset Overview
We evaluate GDF using both real and synthetic datasets to assess its effectiveness in outage prediction and resilience planning. The real dataset records the number of customers affected by outages during the 2018 Nor’easter in Massachusetts [47], while the synthetic outage trajectories are generated using a simplified SIR model.
@@ V-A1 Real Dataset
The real dataset used in this study comprises county-level customer outage counts [13], combined with meteorological measurements and socioeconomic indicators from regions affected by a Nor’easter snowfall event in Massachusetts in 2018 (Fig. 1). Meteorological variables—such as wind speed, temperature, and pressure—are sourced from NOAA’s High-Resolution Rapid Refresh (HRRR) model [14]. Socioeconomic and demographic data are collected from the U.S. Census Bureau’s American Community Survey [15] and include median household income, median age, the number of food stamp recipients, the unemployment rate, the poverty rate, college enrollment, mean travel time to work, and average household size. These variables capture economic and mobility factors that may influence outage recovery dynamics and the effectiveness of emergency response efforts.
A key advantage of using outage data from Massachusetts during the 2018 Nor’easter event [47] is that three consecutive snowstorms impacted the power system within a short 15-day period, during which the local infrastructure remained largely unchanged. This allows us to reasonably assume that the outage patterns from these storms follow the same underlying data distribution, making them suitable for widely used train-test evaluation. Accordingly, we use the first storm for training and the second for testing the effects of different frameworks, while excluding the third storm due to its relatively minor impact. Fig. 1 illustrates the spatiotemporal dynamics of the real dataset.
@@ V-B Experimental Setup
To evaluate the effectiveness of the proposed Global Decision-Focused (GDF) Neural ODE framework, we conduct experiments on two application domains: power line undergrounding [48, 49, 50, 51] and mobile generator deployment as introduced in Section IV. For both problems, predictions for the entire time horizon are made before the occurrence of the hazard, reflecting the practical challenge of collecting data in real time during a developing hazard [52]. Consequently, for both problems, when the hazard surpasses a predefined threshold (1% in our experiments), the ODE model’s initial conditions are set, and decisions are made accordingly.
In the generator deployment problem, we assume a single centralized warehouse as the sole source of generators, with no inter-county transfers. Generators are dispatched from the warehouse to affected counties and must eventually return for refueling, following the setup described in Section V-C1. We also assume a uniform travel time parameter, $\delta_{kk^{\prime}}=\delta_{t},\;\forall(k,k^{\prime})\in\mathcal{E}$, to capture the impact of transportation delays. The online baseline method assumes an observation lag, making greedy observe-then-optimize decisions based on data from previous data, reflecting a realistic delay in data availability for decision makers without predictive insights.
@@
The training process consists of two stages. Initially, the neural ODE model is trained using a standard mean squared error (MSE) loss, providing a predictive model baseline. Optimal decisions based on these baseline predictions serve as a Two-stage reference. In the second stage, the baseline model is then used to initialize end-to-end training with the combined loss function in (6), which integrates both decision-focused and prediction-focused objectives, as described in Algorithm 1.
@@
We compare the proposed model, GDF, with spatio-temporal neural ODE models trained solely on MSE loss (Two-stage) and the ground truth optimal decision baseline (True Optimal). We also include results from an online method for the mobile generator deployment problem. Evaluation is based on three metrics: ($i$) prediction accuracy, measured as the mean squared error between predicted and observed outages; ($ii$) decision loss, represented by the total cost for generator deployment or the total System Average Interruption Duration Index (SAIDI) for grid hardening; and ($iii$) regret, defined as the cost difference between the ground truth optimal decision and the decision based on the model’s predictions. For all experiments on synthetic datasets, reported results and standard deviations are calculated with three different random seeds, which generate varying outage curves and weather conditions.
TABLE I: Out-of-Sample Performance for Generator Distribution Problem with Synthetic Data. Results are averaged over 3 repeated experiments with standard error (SE) in brackets.
@@
Two-stage 9336.0 (975.6) 10459.8 (90.6) 142.9 (90.3) 9336.0 (975.6) 11506.6 (98.1) 474.9 (95.7) 9336.0 (975.6) 12746.2 (196.7) 272.4 (207.2)
GDF 7204.1 (1804.8) 10409.9 (79.7) 93.0 (11.8) 7531.5 (1805.8) 11452.1 (80.3) 420.3 (24.3) 8907.7 (2645.2) 12601.9 (58.1) 128.2 (49.3)
TABLE II: Out-of-Sample Performance of Power Line Undergrounding Problem with Real and Synthetic Data
@@
MSE (×$\times 10^{3})$ SAIDI Regret MSE (×$\times 10^{3})$ SAIDI Regret
True Optimal / 218.2 / / 15.6 /
Two-stage 117.4 328.3 112.1 13.5 16.1 0.5
GDF 165.6 312.7 94.5 13.1 15.6 0
V-C Results
This section presents the results of GDF on mobile generator deployment and power line undergrounding, evaluated in terms of predictive accuracy and decision quality. Results on both synthetic and real datasets demonstrate that GDF improves decision-making compared to conventional Two-stage methods, enabling a more effective response to natural hazards.
V-C1 Mobile Generator Deployment Problem
Table I summarizes the out-of-sample performance for the generator deployment problem on synthetic data under different travel time settings ($\delta_{t}=1,5,10$). The proposed method consistently delivers the best decision quality, as reflected in the lowest regret. Notably, as $\delta_{t}$ increases—representing longer transportation delays—the performance improvement of GDF over the two-stage approach becomes increasingly significant. A similar trend is observed under higher transportation cost factors, as shown in Fig. 4, until transportation costs become so high that the best strategy is effectively to remain idle, at which point the improvement from proactive actions with GDF diminishes.
These findings highlight the importance of proactive scheduling and early interventions when facing longer deployment delays or higher transportation costs. It also demonstrates the advantage of GDF in optimizing resource allocation under more challenging, time-sensitive conditions in the face of extreme hazards that threaten grid resilience. As visualized in Fig. 3, the GDF model slightly overestimate outages in critical regions first hit by outage, enabling the reallocation of additional resources to mitigate potential disruptions. While this adjustment in the forecast is subtle, it leads to a significant reduction in regret. Remarkably, the GDF models also achieve MSE comparable to that of the MSE-trained model in Table I, albeit with higher variance. This indicates that improved decision quality was not achieved at the expense of predictive accuracy but rather through targeted and meaningful adjustments in the forecasts.
Additionally, we conducted extensive experiments on synthetic data for the mobile generator deployment problem under various settings, as shown in Fig. 4. We observed that the performance advantage of GDF over the Two-stage and greedy methods narrows as customer interruption costs, operational costs, or the number of available generators increase. This behavior is expected, as higher system resources or reduced flexibility diminish the need for globally optimized interventions—resulting in less improvement with GDF.
V-C2 Power Line Undergrounding
Table II presents the out-of-sample performance on both the Nor’easter, MA, 2018 event and the synthetic dataset. For the real dataset, the proposed GDF model—despite a slightly higher MSE—delivers improved decision quality, achieving a lower SAIDI and reduced regret compared to the Two-stage baseline. On the synthetic dataset, GDF similarly outperforms the Two-stage method in decision quality while maintaining a low MSE, possibly due to the relatively simple prediction tasks. These results confirm that GDF leads to better overall decision performance in real and synthetic settings.
Overall, the experimental results on both real and synthetic data demonstrate that GDF achieves substantial improvements in decision quality compared to the traditional two-stage approach. Although both models exhibit comparable prediction accuracy, GDF consistently yields lower regret values, highlighting the benefits of proactive scheduling and decision-focused training.

### v3

@@ III-C Model Estimation
The learning of GDF Neural ODEs is carried out through stochastic gradient descent, where the gradient is calculated using a novel algorithm based on differentiable optimization techniques [amos2017optnet, cvx, zhu2022distributionally]. To enable differentiation through the $\arg\min$ operator in (2) embedded in the global-decision-focused loss, we relax the decision variable $\bm{x}$ from a potentially discrete space to a continuous space. At evaluation, we re-solve the original mixed-integer problems (e.g., Eq. (9)–(14) or Supplementary Material Eq. (1)) using Gurobi [gurobi]. All reported cost, SAIDI, and regret values are computed using integer solutions, ensuring full operational validity.
For combinatorial optimization problems (e.g., power line undergroungding), the problem is reformulated as a differentiable quadratic program, and a small quadratic regularization term is added to ensure continuity and strong convexity [wilder2019melding]. Formally, we replace the original objective in (2) with the following:
@@
where $\rho>0$ ensures differentiability. The full training procedure is summarized in Algorithm 1. More implementation details can be found in the Supplementary Material.
To improve training efficiency, we avoid computing GDF gradients at every iteration, since evaluating the decision loss is computationally costly. Instead, we first pretrain the Neural ODE with the standard MSE loss, which quickly guides the model toward a good predictive initialization using inexpensive gradient computations. Starting from this pretrained state, we then apply a few finetuning iterations with the GDF loss, thereby reducing the number of expensive decision-loss evaluations while still aligning the model with global decision-focused objectives. Empirically, our training time remains comparable to that of RNN and LSTM, underscoring both the tractability and scalability of the proposed approach. More details are provided in Section H of the Supplementary Material.
@@ IV Application: Mobile Generator Deployment
Figure 3: A synthetic example of the mobile generator deployment problem for a system with three cities and five generators ($Q_{w}=5$). The $y$-axis represents the number of outaged households. In this example, the uniform travel time $\delta_{t}=10$, transportation cost is set to $c=400$, the customer interruption cost is $\tau=1$, and the operational cost is $\gamma=2$. Four methods are compared on out‑of‑sample data: the proposed GDF framework (regret = $183.47$), a two‑stage approach (regret = $488.45$), an online baseline with observation lag of 5 (regret = $2846.13$), and the optimal solution with the ground-truth. Details of the synthetic data setup are provided in Section V-A.
Our method is broadly applicable to various proactive decision-making problems in grid operations, particularly in response to potential hazard events. This section highlights the adaptability of GDF by applying it to a mobile generator deployment task [8668633, QIN2024752, 9117454], a representative PATOG problem.
@@
Formally, let $Q_{w}\in\mathbb{Z}_{+}$ denote the initial inventory of generators at warehouse $w$. For simplicity, we assume that warehouses also function as staging areas, where generators are initially stored and returned after deployment. We assume that each generator has a fixed maximum capacity, capable of supplying electricity to $N_{g}$ customers, and that a single extreme event is anticipated. The planning horizon is discretized into uniform time periods $\mathcal{T}=\{1,2,\dots,T\}$, where T{T} represents the expected duration of the outage. Let $\mathcal{K}=\{1,2,\dots,K\}$ be the set of $K$ service units (e.g., cities or counties) where generators can be deployed, and $\mathcal{W}=\{1,\dots,W\}$ be the set of warehouses where all generators are initially stored. We denote the stock of generators in unit $k$ at time $t$ as $q_{tk}$.
@@
which consists of three components: ($i$) Transportation Cost: The cost of delivering generators from warehouses to service units. ($ii$) Operation Cost: The fixed operational cost per unit time ($\gamma$) for deployed generators. (iiiiii) Outage Cost: Economic loss due to power outages, determined by the number of customers experiencing outages given the generator deployment decision $\bm{x}$. For simplicity, we define the outage cost as:
@@
where $\tau$ denotes unit customer interruption cost, an adapted version of the Value of Lost Load (VoLL) [6672826], representing the economic impact per household outage per day, and $Y_{k}(t)$ denotes the number of outages in service unit $k$ at time $t$, extracted from the state $\mathbf{S}$ according to (5). This framework primarily evaluates system functionality based on the number of customers experiencing outages ($Y$). It is worth emphasizing that it is flexible to incorporate additional metrics, including the number of restored customers ($R$) and unaffected customers ($U$), if needed.
The transportation and inventory constraints of the mobile generator deployment are as follows:
@@ V-A Dataset Overview
We evaluate GDF using both real and synthetic datasets to assess its effectiveness in outage prediction and resilience planning. The real dataset records the number of customers affected by outages during the 2018 Nor’easter in Massachusetts [wiki:March_11-15_2018_nor'easter], while the synthetic outage trajectories are generated using a simplified SIR model.
@@ V-A1 Real Dataset
The real dataset used in this study comprises county-level customer outage counts [maema2020power], combined with meteorological measurements and socioeconomic indicators from regions affected by a Nor’easter snowfall event in Massachusetts in 2018 (Fig. 1). Meteorological variables—such as wind speed, temperature, and pressure—are sourced from NOAA’s High-Resolution Rapid Refresh (HRRR) model [noaa_hrrr]. Socioeconomic and demographic data are collected from the U.S. Census Bureau’s American Community Survey [USCensusBureau2017ACS] and include median household income, median age, the number of food stamp recipients, the unemployment rate, the poverty rate, college enrollment, mean travel time to work, and average household size. These variables capture economic and mobility factors that may influence outage restoration dynamics and the effectiveness of emergency response efforts.
A key advantage of using outage data from Massachusetts during the 2018 Nor’easter event [wiki:March_11-15_2018_nor'easter] is that three consecutive snowstorms impacted the power system within a short 15-day period, during which the local infrastructure remained largely unchanged. This allows us to reasonably assume that the outage patterns from these storms follow the same underlying data distribution, making them suitable for widely used train-test evaluation. Accordingly, we use the first storm for training and the second for testing the effects of different frameworks, while excluding the third storm due to its relatively minor impact. Fig. 1 illustrates the spatiotemporal dynamics of the real dataset.
@@ V-B Experimental Setup
To evaluate the effectiveness of the proposed Global Decision-Focused (GDF) Neural ODE framework, we design our experiments in two stages: ($i$) testing the predictive accuracy and physical interpretability of the Neural ODE model itself, and ($ii$) applying it to two representative grid resilience planning and operations problems—power line undergrounding [8278121, 7922545, AbiSamra2013, Shea2018] and mobile generator deployment [8668633, QIN2024752, 9117454]. In both applications, the model produces predictions over the full planning horizon before the hazard occurs, reflecting the real-world challenge that utilities cannot rely on real-time data streams during rapidly evolving disasters [9803820]. Once the hazard intensity exceeds a predefined threshold (set to 1% in our experiments), the ODE model’s initial conditions are established, and the corresponding decisions are made.
@@
The training of the Neural ODE follows a two-step process. First, the model is trained with a standard mean squared error (MSE) loss to establish a predictive baseline, from which optimal decisions are derived as a reference point. Building on this initialization, the model is then refined through end-to-end training using the combined loss function in (6), which balances decision-focused and prediction-focused objectives, as detailed in Algorithm 1.
Figure 4: The real outage and restoration trajectories in Indianapolis, IN and their predictions. Left: actual outages and the Neural ODE prediction; the right axis shows restored and unaffected customers for both ground-truth and Neural ODE prediction. The vertical dotted line marks the train–test split. Right: comparison of test-set active-outage predictions from the Neural ODE, RNN (and extended), and LSTM (and extended).
@@
We compare the proposed model, GDF, with spatio-temporal Neural ODE models trained solely on MSE loss (Two-stage) and the ground-truth optimal decision baseline (True Optimal). We also include results from an online method for the mobile generator deployment problem. Evaluation is based on three metrics: ($i$) prediction accuracy, measured as the mean squared error between predicted and observed outages; ($ii$) decision loss, represented by the total cost for generator deployment problem or the total System Average Interruption Duration Index (SAIDI) [ieee1366_2022] for power line undergroudning problem; and (iiiiii) regret, defined as the cost difference between the ground-truth optimal decision and the decision based on the model’s predictions. For all experiments on synthetic datasets, reported results and standard deviations are calculated with three different random seeds, which generate varying outage curves and weather conditions.
TABLE I: Outage prediction performance with Indianapolis Data
Model # Params MSE (Outage) MSE (Restored)
Short Period: 2021‑06‑17 – 2021‑06‑20
RNN 12,673 471.7 N/A
RNN (extended) 12,673 577.2 N/A
LSTM 50,497 455.4 N/A
LSTM (extended) 50,497 567.8 N/A
Neural ODE 2 252.7 3936.0
Long Period: 2023-02-27 – 2023-04-04
RNN 12,673 1106.0 N/A
RNN (extended) 12,673 1502.5 N/A
LSTM 50,497 327.6 N/A
LSTM (extended) 50,497 225.6 N/A
Neural ODE 1,986 93.5 5808.7
  • Notes: RNN/LSTM do not predict $R(t)$; therefore MSE (Restored) is reported as N/A; The test period for the short-period setting corresponds to the last 55 hours and for the long-period setting to the last 96 hours.
TABLE II: Performance of the generator deployment (synthetic data) with varying travel time ($\delta_{t}$)
@@
Neural ODE (Two-stage) 9336.0 (975.6) 10459.8 (90.6) 142.9 (90.3) 9336.0 (975.6) 11506.6 (98.1) 474.9 (95.7) 9336.0 (975.6) 12746.2 (196.7) 272.4 (207.2)
Neural ODE (GDF) 7204.1 (1804.8) 10409.9 (79.7) 93.0 (11.8) 7531.5 (1805.8) 11452.1 (80.3) 420.3 (24.3) 8907.7 (2645.2) 12601.9 (58.1) 128.2 (49.3)
TABLE III: Performance for the generator deployment (synthetic data) with varying number of warehouses ($W$)
Model $W$ = 2 $W$ = 3 $W$ = 4
MSE Cost Regret MSE Cost Regret MSE Cost Regret
True Optimal / 16173.10 0 / 15639.77 (923.76) 0 / 12473.7 (63.8) 0
Online (lag = 1) / 18032.49 (1154.70) 1859.39 (1154.70) / 18032.49 (1154.70) 2392.72 (916.1) / 18032.49 (1154.70) 9026.52 (1154.70)
Online (lag = 3) / 21500.22 (148.37) 5327.12 (148.37) / 21500.22 (148.37) 5860.45 (135.45) / 21500.22 (148.37) 5558.79 (148.37)
RNN (Two-stage) 6356.68 (236.60) 17904.32 (0.58) 1731.32 (0.58) 29537.25 (1304.80) 17637.40 (462.90) 1997.63 (460.86) 29537.25 (1304.80) 18170.73 (460.86) 1997.63 (460.86)
RNN (GDF) 30448.27 (1470.67) 17903.66 (0.00) 1730.55 (0.00) 29840.90 (1279.37) 17637.40 (462.90) 1997.63 (460.86) 29839.47 (1279.97) 18170.73 (460.86) 1997.63 (460.86)
LSTM (Two-stage) 10906.76 (7882.75) 17659.73 (905.46) 1486.63 (905.46) 10906.76 (7882.75) 17719.59 (983.21) 2079.82 (1795.57) 13277.65 (7321.11) 18252.93 (1795.57) 2079.82 (1795.57)
LSTM (GDF) 9777.48 (6663.49) 18139.86 (1821.40) 872.22 (755.34) 13004.73 (8930.83) 18139.86 (1821.40) 2500.09 (1767.27) 9777.48 (6663.49) 18275.56 (1762.44) 2102.46 (1762.44)
Neural ODE (Two-stage) 1278.43 (27.42) 16214.97 (3.70) 40.51 (1.41) 1278.43 (27.42) 15680.28 (922.45) 39.92 (0.59) 1278.43 (27.42) 16213.61 (1.41) 40.51 (1.41)
Neural ODE (GDF) 1047.47 (89.20) 16184.52 (0.64) 11.42 (0.64) 970.80 (32.34) 15648.84 (924.81) 9.07 (1.06) 976.68 (36.06) 16182.01 (1.34) 8.91 (1.34)
V-C Ablation Study on Neural ODE
To evaluate the predictive accuracy and physical interpretability of the proposed Neural ODE in (5), we conduct an ablation study using real-world failure and restoration data from Indianapolis, IN. This dataset contains $169,123$ outage records spanning 2004–2024, where each record provides pole-level failure and restoration timestamps. By aggregating these records to the city level, we construct ground truth trajectories for the failure process $U(t)$ and restoration process $R(t)$, which serve to assess the model’s compartmental interpretation. We benchmark our approach against standard RNN [rumelhart1986learning] and LSTM [hochreiter1997long] models under identical experimental settings, including enhanced variants trained with all available data preceding the standard training start date in that year (denoted as “extended”). Detailed model architectures for the RNN, LSTM, and Neural ODEs are provided in Section D of the Supplementary Material.
The predictive results in Fig. 4 indicate that our model captures a distinct “surge–plateau–restoration” curve that closely mirrors the observed failure–restoration patterns, while also achieving consistently lower MSE than all baselines across both test periods as shown in Table I. Note that the Neural ODE used for short-term forecasting in Table I contains only two trainable parameters—one governing the failure transmission rate and the other the restoration rate. Despite this highly compact parameterization, our model still delivers superior predictive accuracy compared to all baselines, demonstrating that the underlying physical process is effectively captured by our modeling assumptions. This property is especially advantageous when training data are scarce during hazards.
V-D Two Case Studies: Results and Implications
This section presents the results of GDF on mobile generator deployment and power line undergrounding, evaluated in terms of predictive accuracy and regret. All results are evaluated on out-of-sample test sets for both synthetic and real data. For synthetic settings, we report both mean and standard deviation in brackets over three runs. Results on both synthetic and real datasets demonstrate that GDF improves decision-making compared to conventional Two-stage methods, enabling a more effective response to natural hazards.
V-D1 Mobile Generator Deployment Problem
In this problem, we consider $W$ warehouses that store generators. For simplicity, we require each generator to return to a warehouse after completing a task, permitting only transportation between a city and a warehouse, while disallowing direct travel between cities or between warehouses. Furthermore, we impose a stock-balance constraint ensuring that the final number of generators at each warehouse equals to its initial inventory. Because this is a short-term operational problem, we also benchmark against an online method, a widely used greedy strategy in practice, where decisions are made only after damages are observed, reflecting delays that arise from the absence of predictive insights.
Table II reports the out-of-sample performance on synthetic data with a single warehouse $W=1$ under varying travel times, $\delta_{t}=1,5,10$. The results show that our method consistently achieves the lowest regret, and the performance gap between GDF and other baselines widens as $\delta_{t}$ increases. Table III presents the out-of-sample performance on the same dataset while varying the number of warehouses, $W=2,3,4$. As $W$ increases, the problem complexity grows substantially due to the larger set of decision variables. Even under this added complexity, the proposed Neural ODE consistently outperforms all baselines, achieving both stronger predictive accuracy and markedly lower regret.
These results underscore the value of proactive scheduling and early intervention, particularly under longer deployment delays or larger-scale problems. The success of the model can be explained by its behavior shown in Fig. 3, where the GDF model tends to slightly overestimate outages in the most critical regions initially affected, which enables the reallocation of extra resources to prevent severe disruptions. Although this adjustment in the forecast is subtle, it translates into a substantial reduction in regret. Moreover, the GDF models attain MSE levels comparable to those of the MSE-trained baseline (Table II; Table III), albeit with higher variance. This demonstrates that the gains in decision quality arise not from sacrificing predictive accuracy, but from purposeful and targeted refinements in the forecasts.
Fig. 5 presents additional experiments varying other four key hyperparameters: customer interruption cost ($\tau$), transportation cost ($c$), operation cost ($\gamma$), and the number of generators ($Q_{w}$). We find that the performance gap between GDF and the baselines narrows as all parameters increase, except for transportation cost. This behavior is expected, as higher system resources or reduced flexibility diminish the need for globally optimized interventions, resulting in less improvement with GDF.
V-D2 Power Line Undergrounding
We also investigate power line undergrounding [8278121, 7922545, AbiSamra2013, Shea2018] as a long-term planning application of our framework. The goal is to identify an optimal subset of locations for undergrounding to mitigate risks from future hazards. Unlike the mobile generator deployment problem, this task focuses exclusively on deriving undergrounding decisions by prioritizing the most vulnerable areas based on predicted total damage. Further details of the experimental setup are provided in Section B of the Supplementary Material.
Table IV presents the out-of-sample performance on both the Nor’easter, MA, 2018 event and the synthetic dataset. For the real dataset, the proposed GDF model—despite a slightly higher MSE—delivers improved decision quality, achieving the lowest SAIDI and MSE compared to all baselines. On the synthetic dataset, GDF similarly outperforms the Two-stage method in decision quality while maintaining a low MSE. Both results demonstrate that GDF offers substantial gains in decision quality compared with the traditional two-stage approach while predictive accuracy remains comparable across models. This underscores the value of proactive scheduling and global decision-focused training for long-term planning tasks.
TABLE IV: Performance of the Power Line Undergrounding
@@ V-D2 Power Line Undergrounding
MSE ($\times 10^{9})$ SAIDI Regret MSE ($\times 10^{2})$ SAIDI Regret
True Optimal / 0.90 / / 15.6 /
RNN (Two-stage) 11.3 2.20 1.3 29.5 17.2 1.6
RNN (GDF) 11.3 2.20 1.3 29.8 17.2 1.6
LSTM (Two-stage) 11.2 2.20 1.3 109.1 17.2 1.6
LSTM (GDF) 11.3 2.20 1.3 130.0 15.6 0
Neural ODE (Two-stage) 5.4 1.73 0.8 13.5 16.1 0.5
Neural ODE (GDF) 7.1 1.56 0.7 13.1 15.6 0